# Separating subsets from their images

### v1

@@ 4. Proof of C
$X=\left\{(\alpha,\beta,g)\in\Omega\times\Omega\times G\mid\alpha^{g}=\beta\right\}\,.$
For every $\alpha,\beta\in\Omega$, the set of elements $g\in G$ satisfying $\alpha g=\beta$ is a right coset of the stabiliser $G_{\alpha}$. Therefore,
@@
On the other hand, for every $g$, we can choose $(\alpha,\beta)$ to complete the triplet in $|A\cap B^{g}|$ ways. Hence,
@@
where equality is attained if, and only if, $|A\cap A^{g}|=1$ for every nontrivial $g\in G$.
@@
and thus
@@
Now, we shall prove that the lower bound is met if, and only if, $\Omega$ is the point-set of a projective plane and $G$ is a regular group on $\Omega$. Let
@@
Therefore, $G$ is regular on $\Omega$, and hence the design is symmetric. Using for the last time our hypothesis on the intersections in $A^{G}$, $(A^{G},\Omega)$ is a projective plane, and so is its dual $(\Omega,A^{G})$. ∎
@@
$|A|=\frac{1}{2}+\sqrt{q^{2}+q+1-1+\frac{1}{4}}=\frac{1}{2}+\sqrt{\left(q+\frac{1}{2}\right)^{2}}=q+1.$
Furthermore, we observe that the equality can hold if, and only if, such right-hand side is an integer, which in turn happens if, and only if, $n$ is of the form $n=(m|\alpha|)^{2}+(2|G_{\alpha}|-1)m+1$ for some positive integer $m$. When the right-hand side is not an integer, then the lower bound for $\mathbf{m}(G)$ can be improved by taking the ceiling of that value. This makes us pose the following:
Question 17.
@@
5. Proof of E
We start this section by focusing our attention on groups acting regularly. Then, we will generalize our result to families of permutation group with stabilisers of bounded order.
@@ 5. Proof of E
Let $G$ be a regular permutation group of degree $n$ on $\Omega$. As usual, we identify $G$ with $\Omega$, so that the action we are dealing with is the action of $G$ on itself by right multiplication. Let $A$ be a subset of $G$. Suppose that there is a $g\in G$ such that $A\cap Ag$ is empty. Then $g$ is not an element of $AA^{-1}$, and thus $AA^{-1}$ is a proper subset of $G$. On the other hand, if $A\cap Ag$ is nonempty for every $g\in G$, then $AA^{-1}=G$. Hence, $\mathbf{m}(G)$ is the size of the smallest set $A$ such that $AA^{-1}=G$.
As per 6, a subset $A$ of a group $G$ such that $AA^{-1}=G$ is called a difference basis for $G$. A closely related notion appears in additive combinatorics, where Rédei and Rényi introduced in [39] the concept of covering the finite subset $\{0,1,2,\dots,n\}$ of the integers by a difference set AAA-A, with $A\subseteq\mathbb{Z}$. The problem of constructing such sets $A$ of minimal possible size remains a relevant question in additive number theory (see, for instance, [6, 43]).
These ideas also intersect naturally with classical topics in combinatorics and finite geometry. For instance, planar difference sets are difference bases $A\subseteq G$ with the additional property that, for every $g\in G$, there exists a unique pair $(a,b)\in A^{2}$ such that $g=ab^{-1}$. Such sets play a fundamental role in the construction of finite projective planes. In particular, in his seminal work [45], Singer showed that every finite projective plane admits a cyclic collineation group of order equal to the number of its points. Oberve that this fact is directly connected to the sharpness of the lower bound in C.
@@
The cardinality of a minimal difference basis is bounded below by the size of a minimal difference set, and an upper bound of $3\sqrt{|G|}$ is given in [27]. Determining the exact minimal size of a difference basis remains an active area of research, even for specific families of groups (see [2, 3, 4]). Any improvement to these bounds would lead to immediate improvements in the asymptotic estimates computed in E. For example, for any finite Abelian $p$-group $G$ in its regular right action, where $p$ is a prime with $p\geq 11$, the result of [4] implies that
@@
Proof of E.
Let $G$ be a finite group, and suppose that $G$ acts regularly on itself by right multiplication. By Remark 18, we need to build a difference basis $A$ for $G$ of as small size as possible. By [26, Theorem 1], there exist two positive constants $c_{1}$ and $c_{2}$ and two subsets $B$ and $C$ such that
@@
The computation of the limit follows using C for $\mathbf{f}_{C}(n)$ and the inequality just proved for $\mathbf{F}_{\mathcal{C}}(n)$. ∎
Surprisingly, [26, Theorem 1] relies on the fact that every group which is not cyclic of prime order contains a subgroup $H$ whose order exceeds $|G|^{1/2}$ (see [28] for a proof of this fact), which in turn relies on the classification of finite simple groups. In a classification free context, if $G$ have even order, the best result known is that there exists a subgroup $H$ with $|H|\geq|G|^{1/3}$, as the famous result of Brauer and Fawler in [10] states. On the other hand, if $G$ is of odd order, then, by the Feit–Thompson Theorem [19], $G$ is solvable, and solvability is enough to prove the existence of a subgroup of order exceeding the square root of the order of the group.
As the constant $4/\sqrt{3}$ in [26, Theorem 1] does not depend on the classification of finite simple groups, E remains valid for solvable groups. This prompts us to ask the following question.
Question 19.
@@
Actually, since we have an explicit upper bound for $\mathbf{f}_{D}$ and $\mathbf{F}_{\mathcal{D}}$, an immediate computation shows that, if $\mathbf{c}$ is the maximum of the order of a stabiliser in $\mathcal{C}$, then
@@
We do not put any focus on this precise upper bound because, for the following application, the precise value of $\mathbf{c}$ is not known.
Proof of F.
@@
By the positive solution to Sims’s Conjecture in [14], there exists a function $\mathbf{g}$ such that the order of any stabiliser of a permutation group in $\mathcal{C}$ is at most $\mathbf{g}(d)$. Therefore, combining Theorems C and E with Remark 20, we find that there exists a constant $\mathbf{c}(d)>0$ such that, for every
@@
The claims of F immediately follow. ∎
The proof of G follows verbatim that of F, where $\mathcal{C}$ is the family of $G$ which are $2$-arc-transitive groups of automorphisms of some graphs $\Gamma$, whose valency do not exceed $d$, and the positive solution to Sims’s Conjecture is substituted with the positive solution to Weiss’s Conjecture for $2$-arc-transitive graphs obtained in [20, 47, 48, 50, 51].
This result can be actually generalized to a larger class of automorphism groups of graphs. Given a graph $\Gamma$ and a vertex-transitive group of automorphisms $G$, the local group $L$ of the pair $(\Gamma,G)$ is the permutation group that a vertex-stabiliser induces on the neighbourhood of its fixed point. A permutation group $L$ is graph-restrictive if there exists a constant such that, for every pair $(\Gamma,G)$ that witness $L$ as a local group, then the order of a vertex-stabiliser is bounded by this constant. For instance, the property we have used above is that $2$-transitive permutation groups are graph-restrictive. Therefore, our proof actually extends beyond the scope of G. For instance, in [16], it is conjectured that any semiprimitive permutation group is graph-restrictive. For the curious reader, examples of graph-restrictive permutation groups can be found in [21, 22, 33, 34, 46, 49].
6. Proof of I
As stated in the introduction, to discuss the sharpness of H we need to understand $\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b))$, where the wreath product of symmetric groups is endowed with its imprimitive action.
Proof of I.
Let $G=\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, and let $\Sigma$ be its block system. We denote by $B_{1},\dots,B_{b}$ the blocks of $\Sigma$. Preliminary, we show that $\mathbf{m}(G)\leq a+b-1$. Consider any subset of the domain of the form
$A=B_{1}\cup\{\omega_{2}\}\cup\dots\cup\{\omega_{a}\},$
where $\omega_{i}\in B_{i}$ for $2\leq i\leq a$. Note that |A|=a+b1|A|=a+b-1. Observe that, for every permutation $g\in G$, the set $A^{g}$ contains the block $B_{1}^{g}$, and, as $A$ intersects every block in at least one point, $A$ and $A^{g}$ have nonempty intersection. Hence, $A$ is not self-separable for $G$, and the claim follows.
@@ 6. Proof of I
Assume that $b$ is even. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. Our base case is $b=2$. Let $\ell=|A\cap B_{1}|$, and note that $|A\cap B_{2}|\leq a$-$\ell$. Let $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}C_{2}$ be a permutation such that the blocks $B_{1}$ and $B_{2}$ are switched. If $A$ fills one block, any choice of $g$ works. Otherwise, by $a$-transitivity of $\mathrm{Sym}(a)$, we can assure that $A\cap B_{1}\cap(A\cap B_{2})^{g}$ and $A\cap B_{2}\cap(A\cap B_{1})^{g}$ are both empty. Thus, $A$ and $A^{g}$ have trivial intersection, and $A$ is self-separable for $G$.
Now, assume that the property holds for $b$, and we claim that it also holds for $b+2$. If there does not exist a block $B\in\Sigma$, such that $|B\cap A|\geq\lceil(a+1)/2\rceil$, then, in view of Example 5, there there exists an element $g$ of the base group $\mathrm{Sym}(a)^{b}$ such that, for every $B\in\Sigma$, $|A\cap B\cap(A\cap B)^{g}|=0$. Therefore, $A$ is self-separable for $G$.
@@
Assume that $b\geq 5$ is odd and that $a\neq 3$. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. By the restraint on $b$, our base case is $b=5$. We claim that $|A\cap B_{i}|>a/2$ cannot happen for three blocks at once. To fix the notation, we label this three blocks with the indices $1$, 22 and $3$. If that happens, then
@@
(Observe that the last inequality does not hold if $a=3$.) This contradicts the assumption $|A|\leq a+b-2=a+3$. Hence, the number of blocks containing more than $a/2$ points is at most $2$. We need to deal with three cases based on for how many blocks this occurs. If this occurs for no blocks, we can find a permutation $g$ of the base group such that $A\cap A^{g}$ is empty, as we did for the case $b=2$. Suppose now that there exists exactly one block for which its intersection with $A$ contains more than $a/2$ points, and, without loss of generality, we assume that this block is $B_{1}$, and we let $\ell=|B_{i}\cap A|$. We aim to show that then there exists a block $B_{i}$ for which $|B_{i}\cap A|\leq b$-$\ell$. Indeed, by the way of contradiction, if, for every $2\leq i\leq 5$, $|B_{i}\cap A|\geq a-\ell+1$, then
@@
This goes against the cardinality of $A$. Without loss of generality, $|B_{2}\cap A|\leq a$-$\ell$. Hence, reasoning as in the previous base case, we see that there exists an element $g$ such that $g$ stabilizes the blocks $B_{3}$, $B_{4}$ and $B_{5}$, it swaps the blocks $B_{1}$ and $B_{2}$ and $A\cap A^{g}$ is empty. A similar argument can be carried out to deal with the remaining case. Henceforth, we have proved that every subset $A$ containing less that $a+4$ points is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(5)$ (and $a\neq 3$).
The inductive step for $b$ odd (and $a\neq 3$) is verbatim the one we used for $b$ even.
@@
Then, for every $g\in G$, $A^{g}$ and $A$ will intersect in at least one point in $B_{1}$ or $B_{2}$. Hence $A$ is not self-separable. Our objective is to show that every set $A$ whose cardinality is at most $a$ is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$.
@@
In both cases, we find a contradiction with $|A|\leq a$. Using our usual arguments, our claim implies that there exists a permutation $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$ such that $g$ swaps the blocks $B_{1}$ and $B_{2}$, it fixes $B_{3}$ and $A\cap A^{g}$ is empty. Therefore, we established the theorem in this setting.
Assume that $b=3$ and $a$ is even. Let $A$ be a subset of cardinality at most $a+1$. We can deal with the case with every block containing less that $a/2$ points as usual. Hence, without loss of generality, we suppose that $|B_{1}\cap A|=\ell\geq a/2+1$. It cannot be that, for every $i\in{2,3}$, $|B_{2}\cap A|\geq a-\ell+1$, otherwise
@@
[…]
Assume that $b$ is odd and $a=3$. Observe that, in this setting, there exists a set of size $b+1$ which is not self-separable. Indeed, let $A$ be a subset of $a]\times[b]$ such that, for every $i\leq\lceil(b+1)/2\rceil$, $|A\cap B_{i}|=2$, and $A\cap B_{i}$ is empty otherwise. For every permutation $g\in\mathrm{Sym}(3)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, there exists an index $i$ such that
@@
Finally, we argue by induction on $b$. The base case $b=3$ was already taken care of before, while the induction argument is the same we used already two times in this proof.
This last case exhaust the possibilities for the pairs $(a,b)$, and hence it completes the proof of the first part of the statement of I.
@@
7. Sharpness results for D
In this section, we characterize the permutation groups of degree $n$ with $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ – that is, we characterize the permutation groups that meet the upper bound from D.
We divide the discussion into two fundamentally distinct cases: primitive groups (see Section 7.1) and imprimitive groups (see Section 7.2). Namely, we prove the following characterizations.

### v2

@@ 4. Proof of C
$X=\left\{(\alpha,\beta,g)\in\Omega\times\Omega\times G\mid\alpha\in A,\beta\in B,\alpha=\beta^{g}\right\}\,.$
For every $\alpha,\beta\in\Omega$, the set of elements $g\in G$ satisfying $\alpha=\beta^{g}$ is a right coset of the stabiliser $G_{\alpha}$. Therefore,
@@
On the other hand, for every $g$, we can count the number of pairs $(\alpha,\beta)$ such that $\alpha=\beta^{g},$ which is equal to $|A\cap B^{g}|.$ Hence,
@@
where equality is attained if and only if $|A\cap A^{g}|=1$ for every nontrivial $g\in G$.
@@
and thus solving the quadratic inequality gives
@@
Now, we shall prove that the lower bound is met if and only if $\Omega$ is the point-set of a projective plane and $G$ is a regular group on $\Omega$. Let
@@
Therefore, $G$ is regular on $\Omega$, and hence the design is symmetric. Using our hypothesis on the intersections of sets in $A^{G}$, $(A^{G},\Omega)$ is a projective plane, and so is its dual $(\Omega,A^{G})$. ∎
@@
$\alpha(G)=\frac{1}{2}+\sqrt{q^{2}+q+1-1+\frac{1}{4}}=\frac{1}{2}+\sqrt{\left(q+\frac{1}{2}\right)^{2}}=q+1.$
Furthermore, we observe that the equality can hold if and only if such right-hand side is an integer, which in turn happens if and only if $n$ is of the form $n=(m|\alpha|)^{2}+(2|G_{\alpha}|-1)m+1$ for some positive integer $m$. When the right-hand side is not an integer, then the lower bound for $\mathbf{m}(G)$ can be improved by taking the ceiling of that value. This makes us pose the following:
Question 16.
@@
5. Proof of F
We start this section by focusing our attention on groups acting regularly. Then, we will generalize our result on regular groups to families of permutation group with stabilisers of bounded order.
@@ 5. Proof of F
Let $G$ be a regular permutation group of degree $n$ on $\Omega$. As usual, we identify $G$ with $\Omega$, so that the action we are dealing with is the action of $G$ on itself by right multiplication. Let $A$ be a subset of $G$. Suppose that there is a $g\in G$ such that $A\cap Ag$ is empty. Then $g$ is not an element of $A^{-1}A$, and thus $A^{-1}A$ is a proper subset of $G$. On the other hand, if $A\cap Ag$ is nonempty for every $g\in G$, then $A^{-1}A=G$. Therefore, $A$ is not self-separable if and only if $A^{-1}$ is a difference basis (see 5). Hence, $\mathbf{m}(G)$ is the size of the smallest difference basis for $G$.
A notion closely related to that of difference basis appears in additive combinatorics, introduced in [36] by Rédei and Rényi, where they consider subsets $A\subseteq\mathbb{Z}$ such that the difference set AAA-A covers the finite subset $\{0,1,2,\dots,n\}$. The problem of constructing such sets $A$ of minimal possible size remains a relevant question in additive number theory (see, for instance, [6, 41]).
These ideas also intersect naturally with classical topics in combinatorics and finite geometry. For instance, planar difference sets are difference bases $A\subseteq G$ with the additional property that, for every $g\in G$, there exists a unique pair $(a,b)\in A^{2}$ such that $g=ab^{-1}$. Such sets play a fundamental role in the construction of finite projective planes. In particular, in his seminal work [43], Singer showed that every finite projective plane admits a cyclic collineation group of order equal to the number of its points. Observe that this fact is directly connected to the sharpness of the lower bound in C.
@@
The cardinality of a minimal difference basis is bounded below by the size of a minimal difference set, and an upper bound of $3\sqrt{|G|}$ is given in [24]. Determining the exact minimal size of a difference basis remains an active area of research, even for specific families of groups (see [2, 3, 4]). Any improvement to these bounds would lead to immediate improvements in the asymptotic estimates computed in F. For example, for any finite Abelian $p$-group $G$ in its regular right action, where $p$ is a prime with $p\geq 11$, the result of [4] implies that
@@
Proof of F.
Let $G$ be a finite group, and suppose that $G$ acts regularly on itself by right multiplication. By Remark 17, we need to build a difference basis $A$ for $G$ of as small size as possible. By [23, Theorem 1], there exist two positive constants $c_{1}$ and $c_{2}$ and two subsets $B$ and $C$ such that
@@
Now we can compute the limit for $\mathbf{f}_{C}(n)$ using C and the bound for the limit for $\mathbf{F}_{\mathcal{C}}(n)$ follows immediately. ∎
Surprisingly, [23, Theorem 1] relies on the fact that every group which is not cyclic of prime order contains a subgroup $H$ whose order exceeds $|G|^{1/2}$ (see [25] for a proof of this fact), which in turn relies on the classification of finite simple groups. In a classification free context, if $G$ has even order, the best result known is that there exists a subgroup $H$ with $|H|\geq|G|^{1/3}$, as the famous result of Brauer and Fawler in [10] states. On the other hand, if $G$ is of odd order, then, by the Feit–Thompson Theorem [18], $G$ is solvable, and solvability is enough to prove the existence of a subgroup of order exceeding the square root of the order of the group.
We can reproduce the proof of [23, Theorem 1] under the additional assumption that all groups involved are solvable. In this setting the argument no longer relies on the classification of finite simple groups, and it still yields the same constant $4/\sqrt{3}$, which is sharp for the cyclic group $C_{3}$. Hence, F remains valid for solvable groups in the classification free context. This prompts us to ask the following question.
Question 18.
@@
Actually, since we have an explicit upper bound for $\mathbf{f}_{D}$ and $\mathbf{F}_{\mathcal{D}}$, an immediate computation shows that, if $\mathbf{c}$ is the maximum order of a stabiliser in $\mathcal{C}$, then
@@
We do not focus on this precise upper bound because, for the following application, the precise value of $\mathbf{c}$ is not known.
Proof of G.
@@
By the positive solution to Sims’s Conjecture in [14], there exists a function $\mathbf{g}$ such that the order of any stabiliser of a permutation group in $\mathcal{C}$ is at most $\mathbf{g}(d)$. Therefore, combining Theorems C and F with Remark 19, we find that there exists a constant $\mathbf{c}(d)>0$ such that, for every
@@
The claims of G immediately follow. ∎
The proof of H follows verbatim that of G, where $\mathcal{C}$ is the family of $G$ which are $2$-arc-transitive groups of automorphisms of some graphs $\Gamma$, whose valency do not exceed $d$, and the positive solution to Sims’s Conjecture is substituted with the positive solution to Weiss’s Conjecture for $2$-arc-transitive graphs obtained in [19, 45, 46, 48, 49].
This result can be actually generalized to a larger class of automorphism groups of graphs. Given a graph $\Gamma$ and a vertex-transitive group of automorphisms $G$, the local group $L$ of the pair $(\Gamma,G)$ is the permutation group that a vertex-stabiliser induces on the neighbourhood of its fixed point. A permutation group $L$ is graph-restrictive if there exists a constant such that, for every pair $(\Gamma,G)$ that witness $L$ as a local group, then the order of a vertex-stabiliser is bounded by this constant. For instance, the property we have used above is that $2$-transitive permutation groups are graph-restrictive. Therefore, our proof actually extends beyond the scope of H. For instance, in [15], it is conjectured that any semiprimitive permutation group is graph-restrictive. For the curious reader, examples of graph-restrictive permutation groups can be found in [20, 21, 30, 31, 44, 47].
6. Proof of J
As stated in the introduction, to discuss the sharpness of the bounds in I we need to understand $\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b))$, where the wreath product of symmetric groups is endowed with its imprimitive action.
Proof of J.
Let $G=\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, and let $\Sigma$ be its block system. We denote by $B_{1},\dots,B_{b}$ the blocks of $\Sigma$. First, we show that $\mathbf{m}(G)\leq a+b-1$. Consider any subset of the domain of the form
$A=B_{1}\cup\{\omega_{2}\}\cup\dots\cup\{\omega_{b}\},$
where $\omega_{i}\in B_{b}$ for $2\leq i\leq a$. Note that |A|=a+b1|A|=a+b-1. Observe that, for every permutation $g\in G$, the set $A^{g}$ contains the block $B_{1}^{g}$, and, as $A$ intersects every block in at least one point, $A$ and $A^{g}$ have nonempty intersection. Hence, $A$ is not self-separable for $G$, and the claim follows.
@@ 6. Proof of J
Assume that $b$ is even. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. Our base case is $b=2$. Let $\ell=|A\cap B_{1}|$, and note that $|A\cap B_{2}|\leq a$-$\ell$. Now by $a$-transitivity of $\mathrm{Sym}(a),$ there exists $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}C_{2}$ that switches the blocks $B_{1}$ and $B_{2}$ such that $A\cap B_{1}\cap(A\cap B_{2})^{g}$ and $A\cap B_{2}\cap(A\cap B_{1})^{g}$ are both empty. Thus, $A$ and $A^{g}$ have trivial intersection, and $A$ is self-separable for $G$.
Now, assume that the property holds for $b$, and we claim that it also holds for $b+2$. We again have two cases to consider.
If there does not exist a block $B\in\Sigma$, such that $|B\cap A|\geq\lceil(a+1)/2\rceil$, then, there exists an element $g$ of the base group $\mathrm{Sym}(a)^{b}$ such that, for every $B\in\Sigma$, $|A\cap B\cap(A\cap B)^{g}|=0$. Therefore, $A$ is self-separable for $G$.
@@
Assume that $b\geq 5$ is odd and that $a\neq 3$. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. By the restraint on $b$, our base case is $b=5$. We claim that $|A\cap B_{i}|>a/2$ cannot happen for three blocks at once. Without loss of generality, assume we have $|A\cap B_{i}|>a/2$ for $i=1,2$ and $3.$ Then
@@
(Observe that the last inequality does not hold if $a=3$.) This contradicts the assumption that $|A|\leq a+b-2=a+3$. Hence, the number of blocks containing more than $a/2$ points is at most $2$.
There are three cases to consider, corresponding to the number of blocks with $|A\cap B_{i}|>a/2$.
If there are no such blocks, we can find a permutation $g$ of the base group such that $A\cap A^{g}$ is empty, as we did for the case $b=2$. Suppose now that there exists exactly one block for which its intersection with $A$ contains more than $a/2$ points, and, without loss of generality, we assume that this block is $B_{1}$, and we let $\ell=|B_{1}\cap A|$. We aim to show that then there exists a block $B_{i}$ for which $|B_{i}\cap A|\leq b$-$\ell$. Suppose for a contradiction that, for every $2\leq i\leq 5$, $|B_{i}\cap A|\geq a-\ell+1.$ Then
@@
contradicting the cardinality assumption on $A$. So, without loss of generality, $|B_{2}\cap A|\leq a$-$\ell$. Hence, reasoning as in the previous base case, we see that there exists an element $g$ such that $g$ stabilizes the blocks $B_{3}$, $B_{4}$ and $B_{5}$, it swaps the blocks $B_{1}$ and $B_{2}$ and $A\cap A^{g}$ is empty. A similar argument can be carried out to deal with the remaining case. Henceforth, we have proved that every subset $A$ containing less than $a+4$ points is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(5)$ (and $a\neq 3$).
The inductive step for $b$ odd (and $a\neq 3$) is similar to the one we used for $b$ even.
@@
Then, for every $g\in G$, $A^{g}$ and $A$ will intersect $B_{1}$ or $B_{2}$ in at least one point. Hence $A$ is not self-separable. Our objective is to show that every set $A$ whose cardinality is at most $a$ is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$.
@@
In both cases, we find a contradiction to $|A|\leq a$. Using our usual arguments, our claim implies that there exists a permutation $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$ such that $g$ swaps the blocks $B_{1}$ and $B_{2}$, it fixes $B_{3}$ and $A\cap A^{g}$ is empty. Therefore, we established the theorem in this setting.
Assume that $b=3$ and $a$ is even. Let $A$ be a subset of cardinality at most $a+1$. We can deal with the case with every block containing less than $a/2$ points as usual. Hence, without loss of generality, we suppose that $|B_{1}\cap A|=\ell\geq a/2+1$. It cannot be that both $|B_{2}\cap A|\geq a-\ell+1$ and $|B_{2}\cap A|\geq a-\ell+1$, as that would imply that
@@
contradicting the cardinality assumption on $A$. Without loss of generality, we have that $|A\cap B_{2}|\leq a$-$\ell$. Therefore, there exists a permutation $g$ fixing $B_{3}$ and swapping $B_{1}$ and $B_{2}$ such that $A\cap A^{g}$ is trivial. Thus, $A$ is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$, and the desired equality follows.
Assume that $b$ is odd and $a=3$. Observe that, in this setting, there exists a set of size $b+1$ which is not self-separable. Indeed, let $A$ be a subset of $a]\times[b]$ such that, for every $i\leq\lceil(b+1)/2\rceil$, $|A\cap B_{i}|=2$, and $A\cap B_{i}$ is empty otherwise. Now, for every permutation $g\in\mathrm{Sym}(3)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, there exists an index $i$ such that
@@
Finally, we argue by induction on $b$. The base case $b=3$ was already taken care of before, while the induction argument is the same as above.
This last case exhausts the possibilities for the pairs $(a,b)$, and hence it completes the proof of the first part of the statement of J.
@@
7. Proof of D
In this section, we prove D, that is, we characterise permutation groups of degree $n$ with $\mathbf{m}(G)=\lceil(n+1)/2\rceil$, which is the upper bound from E.
We divide the discussion into two fundamentally distinct cases: primitive groups (see Section 7.1) and imprimitive groups (see Section 7.2).
7.1. Primitive groups meeting the upper bound
The characterisation of primitive permutation groups $G$ meeting the upper bound on the parameter $\mathbf{m}(G)$ from E is summarised in the following lemma.